OpenStudy (anonymous):

Chlorofluorocarbons (CFC's) are organic compounds that have been implicated in ozone depletion. When the CFC known as Freon 12 (CCl2F2) is exposed to UV radiation (wavelength in the range of 10 nm to 400 nm, a bit shorter than visible light), a chlorine atom breaks off from the rest of the molecule. Prove that this is possible by calculating the maximum wavelength capable of breaking the C-Cl bond. Average Bond Energy (KJ/mol): H-H: 435 F-F: 155 Cl-Cl: 242 C-C: 347 C=C: 610 CC triple bound: 836

OpenStudy (anonymous):

353e-9

OpenStudy (anonymous):

You have the enthalpies of the C-C bond and the Cl-Cl bond, both of which are homonuclear. To find the heteronuclear bond enthalpy, C-Cl, you have you use the formula Professor Cima showed in the lecture. It is the geometric mean of the two homonuclear enthalpies + 96.3(difference of the two electronegativities)^2 ext, use E = ħc / λ where E = energy per photon ħ = plancks constant = 6.626x10^-34 J s c = speed of light = 3.00x10^8 m/s λ = wavelength.. E = ħc / λ rearranging λ = ħc / E for C-Cl... λ = (6.626x10^-34 J s) x (3.00x10^8 m/s) / (5.63x10^-19 J) = 3.53x10^-7 m = 353x10^-9 m = 353 nm